### v3

# On Lattice Isomorphism Problems for Lattices from LCD Codes over Finite Rings

Abstract.
These days, post-quantum cryptography based on the lattice isomorphism problem has been proposed. Ducas-Gibbons introduced the hull attack, which solves the lattice isomorphism problem for lattices obtained by Construction A from an LCD code over a finite field. Using this attack, they showed that the lattice isomorphism problem for such lattices can be reduced to the lattice isomorphism problem with the trivial lattice $\mathbb{Z}^{n}$ and the graph isomorphism problem. While the previous work by Ducas-Gibbons only considered lattices constructed by a code over a finite field, this paper considers lattices constructed by a code over a finite ring $\mathbb{Z}/k\mathbb{Z}$, which is a more general case. In particular, when $k$ is odd, an odd prime power, or not divisible by $4$, we show that the lattice isomorphism problem can be reduced to the lattice isomorphism problem for $\mathbb{Z}^{n}$ and the graph isomorphism problem.
Key words and phrases:
Post-quantum cryptography, Lattice-based cryptography, Lattice isomorphism problem, Code equivalence problem, Graph isomorphism problem
2020 Mathematics Subject Classification:
Primary 11T71; Secondary 14G50
*Corresponding author
1. Introduction
1.1. Background
Lattice-based cryptosystems that are based on the lattice isomorphism problem (LIP), such as Hawk, are expected to be post-quantum secure [6]. Hawk is a round 2 candidate for additional signature NIST PQC competition, and it generates and verifies signatures efficiently on all devices, including low-end ones [10]. Due to these features, Hawk has attracted attention, and therefore, it is important to analyze the hardness of LIP. Ducas and Gibbons [7] introduce the hull attack, which solves LIP for lattices obtained by Construction A of LCD codes over finite fields, where LCD codes are codes that have no intersection with their dual codes. This property plays an important role in the hull attack, and they show that the LIP for such lattices are reduced to the LIP with the trivial lattice $\mathbb{Z}^{n}$ (called $\mathbb{Z}$LIP) and the graph isomorphism problem. However, the hull attack can only be applied to lattices constructed from codes over finite fields, not to lattices generated by codes over finite rings, such as $\mathbb{Z}/k\mathbb{Z}$. In this paper, we consider LIP constructed from codes not over finite fields, but over finite rings ${\mathbb{Z}}/k{\mathbb{Z}}$.
1.2. Previous work
Ducas-Gibbons [7] defined the $s$-hull of a lattice, and they show the reduction of LIP of lattices obtained from the code over finite fields to $\mathbb{Z}$LIP and the graph isomorphism problem for graphs with vertex size $2n$ using the $s$-hull, where $n$ is the lattice dimension. It is known that $\mathbb{Z}$LIP can be solved in $2^{\frac{0.292n}{2}+o(n)}$ time by the Blockwise Korkine-Zolotarev (BKZ) algorithm [8]. Additionally, the graph isomorphism problem can be solved in quasipolynomial time [2], and the time complexity of this LIP is as long as that of $\mathbb{Z}$LIP. This result also serves as a counterexample to the LIP hardness conjecture [8] using the gap, which is one of the lattice invariants. Currently, as a conjecture on the difficulty of LIP, one involving a new lattice invariant called the hullgap is being considered [7].
1.3. Our results
The results shown by Ducas-Gibbons are only for the lattices obtained by the LCD codes over prime fields with positive characteristic. We generalize the hull attack of Ducas-Gibbons for the codes over rings ${\mathbb{Z}}/k{\mathbb{Z}}$.
  • $\cdot$
    When $k$ is odd, the LIP for lattices obtained by the free and LCD codes can be reduced to the $\mathbb{Z}$}LIP and graph isomorphism problem for graphs with vertex size $2n$ (4.2).
  • $\cdot$
    When $k$ is an odd prime power, we can assume the code property only with LCD (4.10).
  • $\cdot$
    When $k$ is even and not divisible by $4$, the free and LCD codes obtained by LIP can also be reduced to the $\mathbb{Z}$}LIP and graph isomorphism problem for graphs with vertex size $3n$ (4.14).
Ducas and Gibbons provided a counterexample to the conjecture regarding the hardness of LIP proposed by Ducas and van Woerden [8]. This counterexample involves certain $q$-ary lattices when $q$ is a prime. Accordingly, they define a new lattice invariant, called the hullgap. Using the hullgap, they modify the conjecture in Section 6 of [7], under which these lattices are no longer counterexamples. Our results extend the construction of such $q$-ary lattices to the case where $q$ is not divisible by $4$. All of these lattices serve as counterexamples only to the original conjecture in [8], but not to the modified version in [7]. This supports the validity of the revised conjecture and contributes to the security analysis of LIP.
1.4. Technical novelty
To generalize to codes over rings ${\mathbb{Z}}/k{\mathbb{Z}}$, there are two obstacles. One is that we can not use the result of [3] directly. To overcome this, we generalize the result of [3] for free and LCD codes over ${\mathbb{Z}}/k{\mathbb{Z}}$. The other obstacle is that the element $2\in{\mathbb{Z}}/k{\mathbb{Z}}$ is often not a unit. To overcome this, we define the extended signed closure of the code and obtain the result when $k$ is not divisible by $4$. At the beginning of Section 4, we provide a more detailed description of the technical overview.
1.5. Related work
As in LWE-based lattice cryptography, there exist two types of attacks for LIP-based cryptosystems: attacks for general LIP problems and ones for structured LIP problems. For the latter case, we see a series of tremendous progresses recently [13, 4, 1]. However, the attacks have not yet been succeeded to break Hawk, which is the main target for the structured LIP attack. In contrast to them, our attack target is general LIP as is given in [7].
1.6. Outline
This paper is organized as follows. In Section 2, we provide definitions of lattice, code, and graph, and introduce computational problems related to them. In Section 3, we present the previous work on the hull attack by Ducas-Gibbons. In Section 4, we generalize the hull attack for lattices constructed by codes in the finite ring ${\mathbb{Z}}/k{\mathbb{Z}}$ and consider the case when $k$ is odd and when $k$ is even but not divisible by $4$.
2. Preliminary
2.1. Lattices
Let $L$ be a discrete subgroup of an $n$-dimensional vector space over $\mathbb{R}$}. If there exist $\mathbf{v}_{1},\ldots,\mathbf{v}_{m}$ such that
$L\coloneq\left\{\sum_{i=1}^{m}x_{i}\mathbf{v}_{i}\mid x_{i}\in{\mathbb{Z}}\right\},$
then $L$ is called a lattice. The dual of the lattice $L$, denoted by $L^{*}$, is defined as follows:
$L^{*}\coloneq\{\mathbf{x}\in{\mathbb{R}}^{n}\mid\mathbf{x}\cdot\mathbf{y}\in{\mathbb{Z}}\text{ for all $\mathbf{y}\in L$}\}.$
Let $s$ be a real number. Then, we call $L\cap sL^{*}$ the $s$-hull of $L$ and denote it by $H_{s}(L)$.
Definition 2.1 (Lattice Isomorphism).
Lattices $L_{1}$ and $L_{2}$ are isomorphic if there exists an orthonormal matrix $O$ such that $L_{1}=OL_{2}$.
Definition 2.2 (Lattice Isomorphism Problem).
The Decision-Lattice Isomorphism Problem (Decision-LIP) is to determine whether $L_{1}$ and $L_{2}$ are isomorphic, and the Search-Lattice Isomorphism Problem (Search-LIP) is to find the orthonormal matrix $O$ that transforms $L_{1}$ into $L_{2}$. In particular, when $L_{2}=\mathbb{Z}^{n}$, this problem is called $\mathbb{Z}$LIP. In this paper, we focus only on the Search-LIP and simply denote LIP as this second problem.
Remark 2.3.
In general, the Code Equivalence Problem (CEP) and the Graph Isomorphism Problem (GI), which are defined later, also have the two types: the search version and the decision version. In this paper, we mainly focus on the search versions of these problems. Thus, hereinafter, when we refer to LIP, CEP, or GI, we mean the corresponding search problems.
2.2. Codes
Let $R$ be a finite ring and let $R^{n}$ be a free module over $R$. Then, the linear code $C$ over $R$ with length $n$ is a submodule of $R^{n}$. Especially, if $R$ is a finite field of order $q$ and the dimension of $C$ is $m$, then $C$ is called an $n,m]_{q}$ linear code. When a code $C$ over $R$ is denoted as $C=\{cG\mid c\in R^{m}\}$, where $G$ is a matrix over $R$ with size $m\times n$ and no row of $G$ can be represented as a linear combination of the other rows, then $G$ is called the generator matrix of $C$. A dual code of $C$, denoted as $C^{\perp}$, is defined as follows:
$C^{\perp}\coloneq\{\mathbf{x}\in R^{n}\mid\mathbf{x}\cdot\mathbf{c}=0\text{ for all $\mathbf{c}\in C$ }\}.$
$C\cap C^{\perp}$ is called the hull, and if the hull is trivial, meaning the hull is $\{0\}$, then $C$ is called an LCD code (linear complementary dual code).
Definition 2.4 (Code Equivalence).
Codes $C_{1}$ and $C_{2}$ are called equivalent if there exist a diagonal matrix $D$ and a permutation matrix $P$ such that the following holds:
$C_{1}=DPC_{2}.$
In particular, when all the elements of $D$ are $\pm 1$, this equivalence is called signed permutation code equivalence, and if $D$ is the identity matrix, it is called permutation code equivalence.
Definition 2.5 (Code Equivalence Problem).
The code equivalence problem is the problem of determining the equivalence of $C_{1}$ and $C_{2}$, and, when they are equivalent, finding the matrices $D$ and $P$. In particular, when $C_{1}$ and $C_{2}$ are signed permutation code equivalent, this problem is called the signed permutation code equivalence problem (SPEP), and when $C_{1}$ and $C_{2}$ are permutation code equivalent, it is called the permutation code equivalence problem (PEP).
In general, the code equivalence problem of codes over a finite field ${\mathbb{F}}_{q}$ with length $n$ is equivalent to the permutation code equivalence of closure codes, which are obtained from the original codes and are over a finite field ${\mathbb{F}}_{q}$ with length (q1)n(q-1)n [15, 16]. Hereinafter, let ${\mathbb{Z}}_{k}$ denote ${\mathbb{Z}}/k{\mathbb{Z}}$, and we consider the codes over a finite ring ${\mathbb{Z}}_{k}$. One of the methods of obtaining a lattice from codes over ${\mathbb{Z}}_{k}$ is called Construction A. Let $\iota_{k}$ be a function from $\mathbb{Z}$} to ${\mathbb{Z}}_{k}$ such that if an integer $x$ is congruent to $y$ modulo $k$, then $\iota_{k}(x)=y$. Additionally, we extend the domain of $\iota_{k}$ to ${\mathbb{Z}}^{n}$ such that $\iota_{k}(\mathbf{v})=(\iota_{k}(v_{1}),\ldots,\iota_{k}(v_{n}))$, where $\mathbf{v}=(v_{1},\dots,v_{n})$. Then, the lattice $C+k{\mathbb{Z}}^{n}$ obtained from the code $C$ using Construction A is defined as follows [5]:
$C+k{\mathbb{Z}}^{n}\coloneqq\{\mathbf{x}\in{\mathbb{Z}}^{n}\mid\iota_{k}(\mathbf{x})\in C\}.$
2.3. Graphs
A graph $G=(V,E)$ consists of a finite set $V$ of vertices and a set $E$ of edges, where each edge is a 2-element subset of $V$. If there exists a mapping $w$ from $E$ to $\mathbb{R}$, then $(V,E,w)$ is called a weighted graph. The adjacency matrix of $(V,E,w)$ is a $|V|\times|V|$ matrix with entries
$A_{ij}=\begin{cases}w(\{i,j\})&\text{if }\{i,j\}\in E,\\ 0&\text{if }\{i,j\}\notin E.\end{cases}$
Definition 2.6 (Graph Isomorphism).
Let $G_{1}=(V_{1},E_{1},w_{1})$ and $G_{2}=(V_{2},E_{2},w_{2})$ be weighted graphs, and let $A_{1}$ and $A_{2}$ be the adjacency matrices of $G_{1}$ and $G_{2}$, respectively. $G_{1}$ and $G_{2}$ are called isomorphic if there exists a permutation matrix $P$ such that $A_{1}=P^{\top}A_{2}P$ holds.
Definition 2.7 (Graph Isomorphism Problem).
Graph isomorphism problem (GI) involves distinguishing whether two given graphs are isomorphic, and if they are, determining the permutation matrix $P$.
GI is known to be solvable in quasipolynomial time [2]. Additionally, the permutation code equivalence problem of LCD codes over finite fields can be reduced to GI [3].
3. The hull attack proposed by Ducas-Gibbons
Ducas-Gibbons [7] investigates the LIP of lattices obtained by Construction A of an LCD code over a finite field. They show that the LIP can be reduced to the $\mathbb{Z}$}LIP and GI using the $p$-hull of the lattice, $H_{p}(L)$. In this section, we present the flow of this reduction and introduce their lemmas for comparing with our results. Hereinafter, let $p$ be a prime number.
3.1. Outline of the hull attack [7]
[…]
Figure 1. Reductions of computational problems in the hull attack [7]
Figure 1 describes the reductions of computational problems in the hull-attack by Ducas-Gibbons [7].
3.1 implies that the LIP of specific lattices can be reduced to the $\mathbb{Z}$}LIP and SPEP of LCD codes with length $n$. 3.3 and 3.4 suggest that this SPEP can also be reduced to the PEP of LCD codes with length $2n$. Additionally, using 3.5, we can reduce this PEP to GI. As a result, for any $p$, the LIP of lattices obtained by Construction A from LCD codes over ${\mathbb{F}}_{p}$ can be reduced to the $\mathbb{Z}$}LIP and GI (Figure 1). Note that 3.4 treats with the case that $p\neq 2$. But when $p=2$, SPEP is equivalent to PEP, so by the same argument, this reduction also holds.
3.2. Lemmas for the hull attack
We introduce some lemmas and terms that are used in the hull attack.
Lemma 3.1 (Lemma $9$ [7]).
Let $C$ be an $n,k]_{p}$ LCD code. For any $i\in\{1,2\}$, let $O_{i}$ be an orthonormal matrix of size $n$, and let $L_{i}=O_{i}(C+p\mathbb{Z}^{n})$. Then, the LIP of $L_{1}$ and $L_{2}$ can be reduced to the $\mathbb{Z}$-LIP and SPEP of LCD codes with length $n$.
Definition 3.2 (Signed closure code [7]).
Let $C$ be an $n,k]_{p}$ code. The signed closure code of $C$, denoted as $C^{\pm}$, is a $2n,k]_{p}$ code and is defined as follows:
$C^{\pm}\coloneq\{(x_{1},-x_{1},x_{2},-x_{2},\ldots,x_{n},-x_{n})\mid(x_{i})_{1\leq i\leq n}\in C\}$
Lemma 3.3 (Lemma $10$ [7]).
Let $C$ and $C^{\prime}$ be $n,k]_{p}$ codes. Then, $C$ and $C^{\prime}$ are signed permutation code equivalent if and only if $C^{\pm}$ and $C^{\prime\pm}$ are permutation code equivalent.
Lemma 3.4 (Lemma $11$ [7]).
If $p\neq 2$, $C$ is LCD if and only if $C^{\pm}$ is LCD.
Lemma 3.5 (Theorem $5$ [3]).
Let $C_{1}$ and $C_{2}$ be LCD codes over a finite field ${\mathbb{F}}_{p}$, and let $G_{1}$ and $G_{2}$ be the generator matrices of $C_{1}$ and $C_{2}$, respectively. Then, $C_{1}$ and $C_{2}$ are permutation code equivalent if and only if the two symmetric matrices $G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}$ and $G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}$ are graph isomorphic.
3.1 is obtained from the property of the $p$-hull of a lattice, as shown in the following lemma.
Lemma 3.6 (Lemma $4$ [7]).
Let $C$ be an $n,k]_{p}$ code and let $L=C+p{\mathbb{Z}}^{n}$ be a lattice. Then,
$H_{p}(L)=C\cap C^{\perp}+p{\mathbb{Z}}^{n}.$
The proof of 3.3 follows the same argument as in [15], and 3.4 can be proven by considering the inner product of two vectors in the code.
In [3], it was shown that given an LCD code $C$ and its generator matrix $G$, the matrix $G^{\top}(GG^{\top})^{-1}G$ is a projection from ${\mathbb{F}}_{q}^{n}$ to $C$. Using this matrix, the PEP of the LCD code is reduced to GI and obtain 3.5. Note that if $C$ is over a finite field, then $C$ is LCD if and only if $GG^{\top}$ is non-singular. Hence, we can always define $G^{\top}(GG^{\top})^{-1}G$. However, in general, if $C$ is defined over ${\mathbb{Z}}_{k}$, this does not hold. For the case of our finite ring, see 4.6 and its proof.
4. Our results
The previous work of Ducas and Gibbons considers lattices obtained by Construction A from codes over finite fields. In this section, we generalize this result to lattices obtained from codes over the finite ring ${\mathbb{Z}}_{k}$. To generalize for the codes over ${\mathbb{Z}}/k{\mathbb{Z}}$, one of the obstacles is the existence of a projection from ${\mathbb{Z}}^{k}$ onto the code. 4.6 shows that if a code is free and LCD, then $G^{\top}(GG^{\top})^{-1}G$ is such a projection, where $G$ is a generator matrix of a code. Another obstacle is that $2$ is not a unit when $k$ is even. To resolve this, we define the extended signed closure of a code (4.11) and prove the reduction of SPEP of codes to PEP of the extended signed closure of codes when $k$ is not divisible by $4$ (4.12). Moreover, to prove 4.12, we also provide a proof of 4.4 that closely follows the structure of the proof of 3.3, but introduces permutation matrices, which were not used in the original proof. In Section 4.1, we prove that when $k$ is odd, LIP for lattices constructed from codes over ${\mathbb{Z}}_{k}$ can be reduced in a manner similar to that of Ducas-Gibbons (4.2), and we also focus on the special case when $k$ is a power of an odd prime (4.10). In Section 4.2, we prove a similar reduction when $k$ is even and not divisible by $4$ (4.14).
4.1. When $k$ is odd
To generalize the results of Ducas-Gibbons to ${\mathbb{Z}}_{k}$ codes, we need some restrictions on the codes.
Definition 4.1 (Free code).
Let $R$ be a ring and let $C$ be a linear code over $R$ with generator matrix $G$. $C$ is called a free code if all the row vectors of $G$ are linearly independent over $R$.
Note that $C$ is free and LCD if and only if $GG^{\top}$ is non-singular. Additionally, $C$ is always free if $C$ is a code over a finite field. When $k$ is odd, we have the following theorem.
Theorem 4.2.
Let $k$ be an odd integer and let $C$ be a free and LCD code over ${\mathbb{Z}}_{k}$ with length $n$. For any $i\in\{1,2\}$, let $O_{i}$ be an orthonormal matrix of size $n$, and let $L_{i}=O_{i}(C+k{\mathbb{Z}}^{n})$. Then, the LIP of $L_{1}$ and $L_{2}$ can be reduced to $\mathbb{Z}$}LIP and GI with vertex size $2n$.
To prove 4.2, we generalize 3.1 - 3.5 to ${\mathbb{Z}}_{k}$ codes, which are 4.3 - 4.6. In fact, 3.1 and 3.3 can be generalized to ${\mathbb{Z}}_{k}$ codes for all $k$.
Lemma 4.3.
For any integer $k$, let $C$ be an LCD code over ${\mathbb{Z}}_{k}$ with length $n$. For any $i\in\{1,2\}$, let $O_{i}$ be an orthonormal matrix of size $n$, and let $L_{i}=O_{i}(C+k{\mathbb{Z}}^{n})$. Then, the LIP of $L_{1}$ and $L_{2}$ can be reduced to the $\mathbb{Z}$}-LIP and the SPEP for free and LCD codes.
Lemma 4.4.
Two linear codes $C_{1}$ and $C_{2}$ over ${\mathbb{Z}}_{k}$ with length $n$ are signed permutation code equivalent if and only if $C_{1}^{\pm}$ and $C_{2}^{\pm}$ are permutation code equivalent.
On the other hand, we have to restrict $k$ to be odd to generalize 3.4.
Lemma 4.5.
When $k$ is odd, a linear code $C$ over ${\mathbb{Z}}_{k}$ with length $n$ is free and LCD if and only if $C^{\pm}$ is free and LCD.
Furthermore, to generalize 3.5 to ${\mathbb{Z}}_{k}$, the code must be free.
Lemma 4.6.
Let $C_{1}$ and $C_{2}$ be free and LCD codes over ${\mathbb{Z}}_{k}$ with length $n$ and generator matrices $G_{1}$ and $G_{2}$, respectively. Then, $C_{1}$ and $C_{2}$ are permutation code equivalent if and only if $G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}$ and $G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}$ are graph isomorphic.
Note that $k$ can be any integer in 4.6.
Using these lemmas, we can prove 4.2.
Proof of 4.2. From 4.3, the LIP for lattices obtained by Construction A from any LCD code over ${\mathbb{Z}}_{k}$ with length $n$ can be reduced to the $\mathbb{Z}$}LIP and SPEP for LCD codes over ${\mathbb{Z}}_{k}$ with length $n$. From 4.4, this SPEP can be reduced to the PEP for codes over ${\mathbb{Z}}_{k}$ with length $2n$. Additionally, from 4.5, if $k$ is odd, the code obtained by 4.4 preserves the properties of being free and LCD. Finally, from 4.6, the PEP for free and LCD codes can be reduced to the graph isomorphism problem, yielding 4.2. $\square$
4.1.1. Proofs of 4.3 - 4.6
Next, we prove 4.3 - 4.6 using a similar discussion to that of Ducas-Gibbons. Firstly, we generalize 3.6 to a linear code over ${\mathbb{Z}}_{k}$.
Lemma 4.7.
Let $C$ be a linear code over ${\mathbb{Z}}_{k}$ with length $n$, and let $L=C+k{\mathbb{Z}}^{n}$ be the lattice obtained from the Construction A of $C$. Then,
$H_{k}(L)=C\cap C^{\perp}+k{\mathbb{Z}}^{n}.$
Proof.
First, we show
(4.1) $C^{\perp}+k{\mathbb{Z}}^{n}=k(C+k{\mathbb{Z}}^{n})^{*}.$
For all elements $x\in C^{\perp}+k{\mathbb{Z}}^{n}$ and $y\in C+k{\mathbb{Z}}^{n}$, the inner product $x\cdot y$ is in $k{\mathbb{Z}}$. Therefore, $C^{\perp}+k{\mathbb{Z}}^{n}\subset k(C+k{\mathbb{Z}}^{n})^{*}$. Similarly, for all $z\in k(C+k{\mathbb{Z}}^{n})^{*}$ and $y\in C+k{\mathbb{Z}}^{n}$, from the definition of the dual of a lattice, $z\cdot y\in k{\mathbb{Z}}$. In particular, $\iota_{k}(z)\cdot\iota_{k}(y)\in k{\mathbb{Z}}$ and $\iota_{k}(z)\in C^{\perp}$ imply $\iota_{k}(k(C+k{\mathbb{Z}}^{n})^{*})\in C^{\perp}$. Therefore, $C^{\perp}+k{\mathbb{Z}}^{n}\supset k(C+k{\mathbb{Z}}^{n})^{*}$ and we obtain Equation 4.1. Hence,
$H_{k}(L)$ $=(C+k{\mathbb{Z}}^{n})\cap k(C+k{\mathbb{Z}}^{n})^{*}$
$=(C+k{\mathbb{Z}}^{n})\cap(C^{\perp}+k{\mathbb{Z}}^{n})$
$=(C\cap C^{\perp})+k{\mathbb{Z}}^{n}.$
We can prove 4.3 using 4.7.
Proof of 4.3. Firstly, we calculate $k$ from the given lattices, $L_{1}$ and $L_{2}$. Let $G_{i}$ be the generator matrix of $L_{i}$ with size $m\times n$, then the determinant of $G_{i}G_{i}^{\top}$ is $k^{n-m}$. Therefore, we obtain $k$ by computing the determinant of $GG^{\top}$. Next, we consider the $k$-hull of $L_{i}$, denoted as $H_{k}(L_{i})$. From 4.7 and the assumption that $C$ is LCD, we have
$H_{k}(L_{i})$ $=O_{i}(H_{k}(C+k{\mathbb{Z}}^{n}))$
$=O_{i}((C\cap C^{\perp})+k{\mathbb{Z}}^{n})$
$=kO_{i}{\mathbb{Z}}^{n}.$
Hence, if we solve $\mathbb{Z}$}LIP for lattices $H_{k}(L_{i})$, then we obtain the orthonormal matrix $\hat{O}_{i}$, which transforms $H_{k}(L_{i})$ to $k{\mathbb{Z}}^{n}$. After this point, the reduction is exactly the same as the proof of 3.1. Since $\hat{O}_{i}$ is a product of $O_{i}^{-1}$ and an automorphism of ${\mathbb{Z}}^{n}$, there exists a diagonal matrix $D_{i}$, whose entries are all $\pm 1$, and a permutation matrix $P_{i}$ such that $\hat{O}_{i}=D_{i}P_{i}O_{i}^{-1}$. Therefore, by applying $\hat{O}_{i}$ to $L_{i}$, we obtain $\hat{O}_{i}L_{i}=D_{i}P_{i}(C+k{\mathbb{Z}}^{n})=D_{i}P_{i}C+k{\mathbb{Z}}^{n}$. Now, the problem of determining $DP$ such that $D_{1}P_{1}C=DPD_{2}P_{2}C$ is an SPEP. As a whole, since $\hat{O}_{i}$ is obtained by $\mathbb{Z}$}LIP and $DP$ is computed by SPEP, the proof is complete. $\square$
4.4 and 4.5 are also obtained by the same arguments which were given in Ducas-Gibbons [7]. While the proof of 4.4 is essentially the same as that of Ducas and Gibbons, we provide a proof using permutation matrices in preparation for the proof of 4.12.
Proof of 4.4. Let $i\in\{1,2\}$, let $G_{i}$ be the generator matrix of $C_{i}$, and let $T$ be a matrix as follows:
$T=\begin{pmatrix}1&-1&&&&&&\\ &&1&-1&&&&\\ &&&&\ddots&&&\\ &&&&&&1&-1\end{pmatrix}$.
Then, the generator matrix of $C_{i}^{\pm}$, denoted as $G_{i}^{\pm}$, satisfies $G_{i}^{\pm}=G_{i}T$. Now, if $C_{1}$ and $C_{2}$ are signed permutation code equivalent, then there exist a permutation matrix $P$ and a diagonal matrix whose entries are only $\pm 1$ such that $G_{1}=G_{2}PD$. Since the matrix $DT$ is a matrix whose rows are either the rows of $T$ or the rows of $T$ multiplied by 1-1, there exists a permutation matrix $P^{\prime}$ such that $DT=TP^{\prime}$. By a similar discussion, $PT$ is a matrix whose rows are equal to some rows of $T$, which means there exists a permutation matrix $P^{\prime\prime}$ such that $PT=TP^{\prime\prime}$. In fact, the swapping of the $i$-th row and $j$-th row can be represented as the swapping of the (2i1)(2i-1)-th column and (2j1)(2j-1)-th column, and swapping the $2i$-th and $2j$-th columns. Hence, there exist $P^{\prime}$ and $P^{\prime\prime}$ such that
$G_{1}T=G_{2}PDT=G_{2}TP^{\prime\prime}P^{\prime},$
which implies that $C_{1}^{\pm}$ are permutation code equivalent to $C_{2}^{\pm}$.
On the contrary, we assume that $C_{1}^{\pm}$ are permutation code equivalent to $C_{2}^{\pm}$. From the definition of $C_{i}^{\pm}$, the generator matrices are $G_{i}T$, and from their permutation code equivalence, there exists a permutation matrix $P$ such that $G_{1}T=G_{2}TP$. Now, $TP$ is a matrix obtained by permuting the column vectors of $T$. Therefore, let $G_{1}T=(u_{1},\ldots,u_{2n})$ and $G_{2}T=(v_{1},\ldots,v_{2n})$, then there exists a permutation $\sigma$ corresponding to $P$ such that $u_{i}=v_{\sigma(i)}$. Then, from the same discussion of the proof of 3.3, we can assume that there exist $\tau_{i}$ for any $i\in\{1,\ldots,n\}$ such that
$\{\tau_{i}(2i-1),\tau_{i}(2i)\}=\{2k-1,2k\}$
for some $k\in\{1,\ldots,n\}$ and $\sigma=\prod_{i=1}^{n}\tau_{i}.$
Furthermore, if $\tau_{i}(2i-1)=2j-1$, then this permutation can be represented as the swap of the $i$-th row and the $j$-th row in $T$, and if $\tau_{i}(2i-1)=2j$, then this permutation can also be represented as the swap of the $i$-th row and the $j$-th row in $T$, followed by multiplying 1-1 to the $i$-th and $j$-th rows. This implies that there exists a permutation matrix $P^{\prime}$ and a diagonal matrix $D$ with all entries being $\pm 1$ such that $TP=P^{\prime}DT$, where $P$ corresponds to the permutation $\sigma$. Therefore, $G_{1}T=G_{2}TP=G_{2}P^{\prime}DT$.
Additionally, let $T^{\prime}$ be the matrix as follows:
$T^{\prime}=\begin{pmatrix}1&&&\\ 0&&&\\ &1&&\\ &0&&\\ &&\ddots&\\ &&&1\\ &&&0\end{pmatrix}$.
Then, since $TT^{\prime}$ is the identity matrix, $G_{1}=G_{2}P^{\prime}D$, and we obtain that $C_{1}$ and $C_{2}$ are signed permutation code equivalent. $\square$
Proof of 4.5. The freeness of $C^{\pm}$ is obtained from the fact that the generator matrix of this code is $GT$. For any vector $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})$ in $C$, both
$\mathbf{x}^{\pm}$ $=(x_{1},-x_{1},\ldots,x_{n},-x_{n}),$
$\mathbf{y}^{\pm}$ $=(y_{1},-y_{1},\ldots,y_{n},-y_{n})$
are elements of $C^{\pm}$, and
$\mathbf{x}^{\pm}\cdot\mathbf{y}^{\pm}=2\mathbf{x}\cdot\mathbf{y}.$
Here, since $k$ is an odd integer, $2$ is a unit over ${\mathbb{Z}}_{k}$. Hence, $\mathbf{x}\cdot\mathbf{y}=0$ is equivalent to $\mathbf{x}^{\pm}\cdot\mathbf{y}^{\pm}=0$. From the above, the LCD property of $C^{\pm}$ is equivalent to that of $C$. $\square$
Proof of 4.6. Let $C$ be a free and LCD code over ${\mathbb{Z}}_{k}$ with length $n$, and let $G$ be a generator matrix of $C$ with size $m\times n$. We show that $G^{\top}(GG^{\top})^{-1}G$ is a projection of ${\mathbb{Z}}_{k}^{n}$ onto $C$. Since $C$ is LCD , $(C\cap C^{\perp})^{\perp}={\mathbb{Z}}_{k}^{n}$. Additionally, since $C$ is free, the size of the generator matrix of $C^{\perp}$, denoted as $H$, is $(n-m)\times n$. From Theorem 3.1 [11], the generator matrix of ${\mathbb{Z}}_{k}^{n}=(C\cap C^{\perp})^{\perp}=C\oplus C^{\perp}$ is given by
$\begin{pmatrix}G\\ H\end{pmatrix}$.
Since $GG^{\top}$ is a non-singular matrix, we can define the following matrix:
$\begin{pmatrix}G^{\top}(GG^{\top})^{-1}&H^{\top}(HH^{\top})^{-1}\end{pmatrix}$,
which is the inverse matrix of $\begin{pmatrix}G\\ H\end{pmatrix}$. Let $\sigma_{C}$ and $\sigma_{C^{\perp}}$ be the projections onto $C$ and $C^{\perp}$, respectively, which satisfy
$\sigma_{C}\circ\sigma_{C^{\perp}}=\sigma_{C^{\perp}}\circ\sigma_{C}=0,\quad\sigma_{C}(\mathbf{v})+\sigma_{C^{\perp}}(\mathbf{v})=\mathbf{v}.$
Then, there exist $\mathbf{x}_{C}\in{\mathbb{Z}}_{k}^{m}$ and $\mathbf{x}_{C^{\perp}}\in{\mathbb{Z}}_{k}^{n-m}$ such that $\sigma_{C}(\mathbf{v})=\mathbf{x}_{C}G$ and $\sigma_{C^{\perp}}(\mathbf{v})=\mathbf{x}_{C^{\perp}}H$. This implies
$\mathbf{v}$ $=\sigma_{C}(\mathbf{v})+\sigma_{C^{\perp}}(\mathbf{v})$
$=\mathbf{x}_{C}G+\mathbf{x}_{C^{\perp}}H$
$=\mathbf{x}\begin{pmatrix}G\\ H\end{pmatrix}$
Therefore,
$\mathbf{x}=\mathbf{v}\begin{pmatrix}G^{\top}(GG^{\top})^{-1}&H^{\top}(HH^{\top})^{-1}\end{pmatrix}$
and
$\sigma_{C}(\mathbf{v})$ $=\mathbf{v}G^{\top}(GG^{\top})^{-1}G,$
$\sigma_{C^{\perp}}(\mathbf{v})$ $=\mathbf{v}H^{\top}(HH^{\top})^{-1}H.$
From the above, we have
$C=\text{Im}(\sigma_{C})=\left\{\mathbf{v}G^{\top}(GG^{\top})^{-1}G\mid\mathbf{v}\in{\mathbb{Z}}_{k}^{n}\right\}.$
If $C_{1}$ and $C_{2}$, which are free and LCD over ${\mathbb{Z}}_{k}$, are permutation code equivalent, then there exists $P$ such that $G_{1}=G_{2}P$. Then,
$G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}$ $=(G_{2}P)^{\top}(G_{2}P(G_{2}P)^{\top})^{-1}G_{2}P$
$=P^{\top}G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}P,$
and we obtain that $G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}$ and $G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}$ are graph isomorphisms. On the contrary, assume there exists $P$ such that
$G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}=P^{\top}G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}P.$
Then,
$C_{1}$ $=\{\mathbf{v}G_{1}^{\top}(G_{1}G_{1}^{\top})^{-1}G_{1}\mid\mathbf{v}\in{\mathbb{Z}}_{k}^{n}\}$
$=\{\mathbf{v}P^{\top}G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}P\mid\mathbf{v}\in{\mathbb{Z}}_{k}^{n}\}$
$=\{\mathbf{v}G_{2}^{\top}(G_{2}G_{2}^{\top})^{-1}G_{2}P\mid\mathbf{v}\in{\mathbb{Z}}_{k}^{n}\}$
$=C_{2}P$
and $C_{1}$ is permutation code equivalent to $C_{2}$. $\square$
Note that the most important part of the proof of 4.6 is that $GG^{\top}$ is non-singular. Since codes over any field are always free, we can prove 3.5 for codes assuming only that they are LCD.
4.1.2. The specific case when $k$ is an odd prime power
In fact, when $k$ is an odd prime power, we can eliminate the property of freeness of the code. By the following argument, if a ring $R$ has the property called FCR (finite chain ring), then the LCD code over $R$ is always free.
Definition 4.8 (Finite chain ring (FCR)).
A finite commutative ring with unity is called a finite chain ring (FCR) if, for any two ideals, one is contained in the other.
Proposition 4.9 ([7]).
Let $R$ be an FCR, let $C$ be a code over $R$ with length $n$, and let $G$ be a generator matrix of $C$. Then, the following are equivalent:
  • (i)
    $C$ is LCD.
  • (ii)
    $GG^{\top}$ is non-singular.
Note that $GG^{\top}$ is non-singular if and only if $C$ is free and LCD code. Liu-Liu [12] shows that 4.9 holds for specific LCD codes, and Durǧun [9] shows 4.9 holds for any LCD code. It is also known that ${\mathbb{Z}}_{k}$ is an FCR if $k$ is a prime power, so in this case, we can obtain 4.10.
Corollary 4.10.
Let $q$ be an odd prime power, and let $C$ be an LCD code over $\mathbb{Z}_{q}$ with length $n$. For any $i\in\{1,2\}$, let $O_{i}$ be an orthonormal matrix of size $n$, and let $L_{i}=O_{i}(C+p^{a}\mathbb{Z}^{n})$. Then, LIP for $L_{1}$ and $L_{2}$ can be reduced to $\mathbb{Z}$LIP and GI with $2n$ vertices.
4.2. The case $k$ is even and not divisible by 44
Additionally, we consider the case when $k$ is even. In this case, the obstacle to generalization is 4.5. To resolve this, we modify the signed closure code.
Definition 4.11 (Extended signed closure).
Let $k=2m$ be an even number, where $m$ is odd, and let $T_{m}$ be a matrix of size $n\times 3n$ as follows:
$T_{m}=\begin{pmatrix}T&mI_{n}\end{pmatrix}$,
where $T$ is defined in 4.4. For any code $C$ over ${\mathbb{Z}}_{k}$ with length $n$, the extended signed closure code of $C$, denoted as $C^{e}$, is a code over ${\mathbb{Z}}_{k}$ with length $3n$, defined as follows:
$C^{e}\coloneq\{cT_{m}\mid c\in C\}.$
Extended signed closure code is a code whose first $2n$ rows are equivalent to a signed closure code, and after the $(2n+1)$-th row, all entries are $m$ or 0. When $k$ is even and not divisible by $4$, the SPEP for free and LCD codes over ${\mathbb{Z}}_{k}$ with length $n$ can be reduced to the PEP for free and LCD codes over ${\mathbb{Z}}_{k}$ with length $3n$ using the extended signed closure code. Firstly, we show 4.12, which corresponds to the generalization of 4.4.
Lemma 4.12.
Let $m$ be an odd integer and $k=2m$. The signed permutation code equivalence of $C_{1}$ and $C_{2}$, which are over ${\mathbb{Z}}_{k}$, is equivalent to the permutation code equivalence of $C_{1}^{e}$ and $C_{2}^{e}$.
Proof.
In this proof, if we denote the column vectors $\{\mathbf{v}_{1},\ldots,\mathbf{v}_{r}\}$, we consider this as a multiset and
$\{\mathbf{u}_{1},\ldots,\mathbf{u}_{r}\}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{r}\}$
denote this as an equivalence of multisets. Let $n$ be the length of $C_{1}$ and $C_{2}$. For any $i\in\{1,2\}$, let $G_{i}$ be a generator matrix of $C_{i}$, and let $G_{i}^{e}$ be a generator matrix of $C_{i}^{e}$. Then, from the definition, $G_{i}^{e}=G_{i}T_{m}$. Firstly, we assume that $C_{1}$ and $C_{2}$ are signed permutation code equivalent, i.e., there exists a permutation matrix $P$ and a diagonal matrix $D$ such that $G_{1}=G_{2}PD$. Thus, $G_{1}^{e}=G_{1}T_{m}=G_{2}PDT_{m}$. Since m=mm=-m,
$DT_{m}=\begin{pmatrix}DT&DmI_{n}\end{pmatrix}=\begin{pmatrix}DT&mI_{n}\end{pmatrix}$.
From 4.4, since there exists a permutation matrix $P^{\prime}$ such that $DT=TP^{\prime}$, there also exists a permutation matrix $P^{\prime}_{3n}$ such that $P^{\prime}_{3n}=\begin{pmatrix}P^{\prime}&0\\ 0&I_{n}\end{pmatrix}$ and $DT_{m}=T_{m}P^{\prime}$. From the proof of 4.4, because there exists a permutation matrix $P^{\prime\prime}$ such that $PT=TP^{\prime\prime}$, $PT_{m}=\begin{pmatrix}PT&Pm\end{pmatrix}=\begin{pmatrix}TP^{\prime\prime}&mI_{n}P\end{pmatrix}=T_{m}\begin{pmatrix}P^{\prime\prime}&0\\ 0&P\end{pmatrix}$. Therefore,
$G_{1}^{e}=G_{2}PDT_{m}=G_{2}T_{m}P_{3n}^{\prime}P^{\prime\prime}=G_{2}^{e}P^{\prime\prime\prime},$
this implies that $C_{1}^{e}$ is permutation code equivalent to $C_{2}^{e}$. On the contrary, we assume there exists a permutation matrix such that $G_{1}T_{m}=G_{2}T_{m}P$. Let
$G_{1}T_{m}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{3n}),\quad G_{2}T_{m}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{3n}),$
and let $\sigma$ be a permutation corresponding to $P$, such that $\mathbf{u}_{i}=\mathbf{v}_{\sigma(i)}$, and define $U_{m}$ and $V_{m}$ as follows:
$U$ $\coloneq\{\mathbf{u}_{1},\cdots,\mathbf{u}_{3n}\},$
$V$ $\coloneq\{\mathbf{v}_{1},\cdots,\mathbf{v}_{3n}\},$
$U_{m}$ $\coloneq\{\mathbf{u}_{i}\in U\mid\mathbf{u}_{i}=m\mathbf{u}_{i}\},$
$V_{m}$ $\coloneq\{\mathbf{v}_{i}\in V\mid\mathbf{v}_{i}=m\mathbf{v}_{i}\}.$
Note that all entries of the column vector $\mathbf{u}_{i}$ are 0 or $m$ if and only if $\mathbf{u}_{i}\in U_{m}$, and $V_{m}$ is the same. Especially, if $i\geq 2n+1$, then $\mathbf{u}_{i}\in U_{m}$ and $\mathbf{v}_{i}\in V_{m}$.
Since $U=V$, $V_{m}=U_{m}$, and $\sigma$ can be represented as the product of permutations $\sigma_{m}$, which maps $V_{m}$ to $U_{m}$, and a permutation $\tau$, which maps $V\setminus V_{m}$ to $U\setminus U_{m}$. From now on, we show the existence of $P^{\prime}$ such that $G_{1}T_{m}=G_{2}T_{m}P^{\prime}$ and $P^{\prime}=\begin{pmatrix}P_{2n}&0\\ 0&P_{n}\end{pmatrix}$, where $P_{2n}$ and $P_{n}$ are permutation matrices of size $2n$ and $n$, respectively. Note that $P^{\prime}$ represents the product of permutations $\sigma_{1}$ and $\sigma_{2}$, where $\sigma_{1}$ is a permutation of the first $2n$ rows and $\sigma_{2}$ is a permutation of the last $n$ rows. Define
$U_{\leq 2n}$ $\coloneq\{\mathbf{u}_{i}\in U_{m}\mid i\leq 2n\},$
$V_{\leq 2n}$ $\coloneq\{\mathbf{v}_{i}\in V_{m}\mid i\leq 2n\}.$
Then, if $U_{\leq 2n}=V_{\leq 2n}$ then such a $P^{\prime}$ exists. Define $V_{m}^{c}\coloneq V\setminus V_{m}$ and $U_{m}^{c}\coloneq U\setminus U_{m}$. Then, $\tau(V_{m}^{c})=U_{m}^{c}$. From the definition of the extended signed closure code, if $\mathbf{u}_{i}\in U_{m}^{c}$, then $\mathbf{u}_{\lceil\frac{i}{2}\rceil+2n}=m\mathbf{u}_{i}$. Define
$U_{m}^{c+}$ $\coloneq\left\{\mathbf{u}_{\frac{i}{2}+2n}\mid\mathbf{u}_{i}\in U_{m}^{c},\text{ $i$ is even}\right\},$
$U_{\leq 2n}^{+}$ $\coloneq\left\{\mathbf{u}_{\frac{i}{2}+2n}\mid\mathbf{u}_{i}\in U_{\leq 2n},\text{ $i$ is even}\right\},$
$V_{m}^{c+}$ $\coloneq\left\{\mathbf{v}_{\frac{i}{2}+2n}\mid\mathbf{v}_{i}\in V_{m}^{c},\text{ $i$ is even}\right\},$
$V_{\leq 2n}^{+}$ $\coloneq\left\{\mathbf{v}_{\frac{i}{2}+2n}\mid\mathbf{v}_{i}\in V_{\leq 2n},\text{ $i$ is even}\right\}.$
Since $U_{m}^{c}=V_{m}^{c}$, $U_{m}^{c+}=V_{m}^{c+}$. Because $U_{m}=U_{m}^{c+}\cup U_{\leq 2n}\cup U_{\leq 2n}^{+}$ and $V_{m}$ is the same, we obtain
$U_{\leq 2n}\cup U_{\leq 2n}^{+}=V_{\leq 2n}\cup V_{\leq 2n}^{+}.$
Additionally, since $m=-m=m^{2}$, we have
$\{\mathbf{u}_{i}\in U_{\leq 2n}\mid\text{$i$ is even}\}$ $=\{\mathbf{u}_{i}\in U_{\leq 2n}\mid\text{$i$ is odd}\}$
$=\{\mathbf{u}_{i}\in U_{\leq 2n}^{+}\},$
which shows that the multiset $\{\mathbf{u}_{i}\in U_{\leq 2n}\cup U_{\leq 2n}^{+}\}$ contains all vectors of $\{\mathbf{u}_{i}\in U_{\leq 2n},\text{$i$ is odd}\}$ exactly three times. The same equation holds for $V_{\leq 2n}$. Therefore,
$U_{\leq 2n}=V_{\leq 2n}.$
From the above, there exists a permutation matrix $P^{\prime}=\begin{pmatrix}P_{2n}&0\\ 0&P_{n}\end{pmatrix}$ such that
$\begin{pmatrix}G_{1}T&mG_{1}\end{pmatrix}$ $=G_{2}T_{m}P^{\prime}$
$=\begin{pmatrix}G_{2}TP_{2n}&mG_{2}P_{n}\end{pmatrix}$.
This especially implies $G_{1}T=G_{2}TP_{2n}$ and $mG_{1}=mG_{2}P_{n}$. The first equation shows a permutation code equivalence of $C_{1}^{\pm}$ and $C_{2}^{\pm}$, and from 4.4, $C_{1}$ and $C_{2}$ are signed permutation code equivalent. Therefore, if $C_{1}^{e}$ and $C_{2}^{e}$ are permutation code equivalent, then $C_{1}$ and $C_{2}$ are signed permutation code equivalent. ∎
Lemma 4.13.
Let $m$ be an odd integer and $k=2m$. A code $C$ over ${\mathbb{Z}}_{k}$ is free and LCD if and only if $C^{e}$ is free and LCD.
Proof.
The freeness of $C^{e}$ also holds by the same argument as in 4.5. We now consider the LCD property. For any vectors $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})$ in $C$,
$\mathbf{x}^{e}$ $=(x_{1},-x_{1},\ldots,x_{n},-x_{n},mx_{1},\cdots,mx_{n}),$
$\mathbf{y}^{e}$ $=(y_{1},-y_{1},\ldots,y_{n},-y_{n},my_{1},\cdots,my_{n})$
are also in $C^{e}$, and
$\mathbf{x}^{e}\cdot\mathbf{y}^{e}=(m^{2}+2)\mathbf{x}\cdot\mathbf{y}.$
Since $k=2m$ and $m$ is odd, $m^{2}+2$ is a unit over ${\mathbb{Z}}_{k}$. Therefore, $\mathbf{x}\cdot\mathbf{y}=0$ and $\mathbf{x}^{e}\cdot\mathbf{y}^{e}=0$ are equivalent. This implies that $C$ is LCD if and only if $C^{e}$ is LCD. ∎
From 4.12 and 4.13, we can generalize 4.2 for $k$ being an even number and not divisible by $4$.
Theorem 4.14.
Let $k$ be an even number and not divisible by $4$, and let $C$ be a free and LCD code over ${\mathbb{Z}}_{k}$ with length $n$. For any $i\in\{1,2\}$, let $O_{i}$ be an orthonormal matrix of size $n$ and define $L_{i}=O_{i}(C+k{\mathbb{Z}}^{n})$. Then, the LIP of $L_{1}$ and $L_{2}$ can be reduced to $\mathbb{Z}$}LIP and GI for graphs with vertex size $3n$.
Proof.
From 4.3, the LIP for lattices obtained by Construction A from any LCD code over ${\mathbb{Z}}_{k}$ with length $n$ can be reduced to the $\mathbb{Z}$}LIP and SPEP for LCD codes over ${\mathbb{Z}}_{k}$ with length $n$, and from 4.12, this SPEP can be reduced to the PEP for codes over ${\mathbb{Z}}_{k}$ with length $3n$. Additionally, from 4.13, the code obtained by 4.12 preserves the properties of being free and LCD. Finally, from 4.6, the PEP for free and LCD codes can be reduced to the graph isomorphism problem, yielding 4.14. ∎
Acknowledgement
This work is supported by JSPS Grant-in-Aid for Scientific Research(C) JP22K11912, JST K Program Grant Number JPMJKP24U2, and MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) JPMXS0120319794.